OpenStudy (anonymous):

show that (3 sin A - cos 2A-1) / (sin² A+2 sin A) = 2 - cosec A

OpenStudy (anonymous):

rearrange the right side to look more attainable \[\frac{ 3sinA-\cos2A-1 }{ \sin^{2}A+2sinA }=2-\frac{ 1 }{ sinA }\] redoing the right again; left: remember your double angles to make cos2A = 1-sin^2A \[\frac{2sin^2A+ 3sinA-2 }{ \sin^{2}A+2sinA }=\frac{ 2sinA-1 }{ sinA }\] factor top and bottom and simplify